Define equilibrium constant and for the following reaction


2NH3(g) ⇌ 2N2(g) +3H2 (g)


i. Write the expression for Kc.


ii. Relate Kc to Kp, and


iii. Relate Kc


to Kx


1
Expert's answer
2022-04-06T07:01:01-0400

Kc​=[N2​][H2​]3
[NH3​]2​,Kp​=pN2​×pH2​3​
pNH3​2​​Δn=2−(1+3)=−2∴Kp​=Kc​(RT)Δn=Kc​(RT)−2=Kc​/(RT)2